OpenStudy (anonymous):

lim x-> -2 from the right (x+3) |x+2|/(x+2)

OpenStudy (anonymous):

if \(x<-2\) you have \(|x+2|=-x-2\) and so \[\frac{|x+2|}{x+2}=\frac{-x-2}{x+2}=-1\]

OpenStudy (anonymous):

so if \(x<-2\) your expression is \((x+3)\times (-1)=-x-3\)

OpenStudy (anonymous):

isn't that 0/0?

OpenStudy (anonymous):

oh nevermind...thank you!

OpenStudy (anonymous):

yw